# Information Flow Routes: Automatically Interpreting Language Models at Scale

### v1

@@ 3.2 Greater-than
For the overall contributions in the greater-than task (Figure 6, left), we also see that (i) task-specific heads discovered via patching are among the most important heads in the information flow routes, and (ii) many other heads are important. As before, let us now look at the differences with the contrastive templates.
@@ 3.3 OPT-125m vs GPT2-Small
Additionally, we experiment with OPT-125m which has the same number of layers and heads than GPT2-small. We find that the information flow routes for IOI and greater-than tasks are similar for OPT-125M and GPT2-small (Appendix B).
@@ 3.4 Hundred Times Faster than Patching
Obtaining information flow routes involves a two-step process. First, we run a forward pass and cache internal activations. Then, we obtain edge importances to build the subgraph as depicted in Algorithm 3. For comparison purposes, we contrast our approach with the ACDC algorithm introduced by (Conmy et al., 2023), an automated circuit discovery algorithm based on patching. According to their findings, on the IOI task with a batch of 50 examples, it requires 8 minutes on a single GPU to discover the circuit. In contrast, our method accomplishes this task in 5 seconds, around 100x in time reduction.
@@ 4 General Experiments
Information flows inside the network via mechanisms implemented in the model. In this section, we look at general patterns in component importances, and try to understand if important mechanisms depend on the type of tokens processed.
Setting.
Using the full information flow graph (Figure 2) of Llama 2-7B, we look at all the immediate connections with each residual stream. We then record the importance values of all the sub-edges corresponding to individual attention heads (i.e., $e_{i,j}^{h}$), as well as FFN blocks. We use a subset of 1000 sentences from the C4 dataset (Raffel et al., 2020).
@@ 4.1 Component Importance for POS
First, let us see how component importance depends on parts of speech (POS). For this, we pack per-prediction component importances into vectors [^4] Each vector corresponding to the $i$-th position is defined as $\left(\sum_{j}e^{1,1}_{i,j},\sum_{j}e^{1,2}_{i,j},\ldots,\sum_{j}e^{L,H}_{i,j},e^{\mbox{ffn}_{1}}_{i},\ldots,e^{\mbox{ffn}_{L}}_{i}\right)$. and apply t-SNE (van der Maaten & Hinton, 2008). Figure 8 shows the resulting projection with datapoints colored according to either input or next token [^5] Here, we take the next token from the dataset and not the one generated by the model. While this adds some noise to the results, we expect at least parts of speech of the reference and the predicted tokens to be similar. part of speech. [^6] We obtain part of speech tags using NLTK library (universal tagset, Petrov et al. (2012)) at the word level and assign a tag to all subwords of the word.
@@
Looking in a bit more detail at the input tokens, Figure 8c shows that contribution patterns strongly depend on whether the current token is the first or a later subword of some word. Clearly, there are some model components specialized for first or later subwords – we confirm this later in Section 4.3. Note also that Figure 8c explains the two distinct clusters for numbers we saw in Figure 8a (purple): they separate number tokens into first and later digits.
@@
Finally, Figure 8b shows the same datapoint colored by part of speech of the next token. Comparing this to Figure 8a, we see that contribution patterns depend more on input tokens rather than output tokens. This might be because in the lower part of the network, a model processes inputs in a generic manner, while the higher network part (that prepares to generate output) is more fine-grained and depends on more attributes than part of speech.
[…]
@@ 4.3 Positional and Subword Merging Heads
Some of our observations above hint at certain attention head functions that might be overall important. First, in Section 3.1 we showed that previous tokens heads are generally important in the IOI task, both for target examples and the contrastive baseline – this attention head function might be important in general. Next, in Section 4.1 we noticed clear difference between component contributions for first and later subwords – this suggests that some model components might be responsible for that.
In what follows, we are going to look at two attention head functions found earlier for machine translation, positional and subword merging heads, and see whether they are generally important. Differently from previous work, we define a head function based on token contributions within head and not attention weights as done previously. This is reasonable since attention weights might not reflect influences properly (Bastings & Filippova (2020); Kobayashi et al. (2020), among others).
@@
[…]
We see subword merging heads in the bottom part of the network (Figure 9a, red), and are among the most important heads in the corresponding layers. This is interesting: previous work that noticed such head roles did not study their general relevance (Correia et al., 2019). Note that these subword merging heads are important for later subwords and not important otherwise – this explains the clusters we saw in Figure 8c.
We would like to highlight that for language models, we are the first to talk about general importance of attention heads and to find that some of the most important heads are previous token and subword merging heads. Future work might explore the functions of other important heads in the model.
@@ 4.4 Peculiar Information Flow Patterns, or Periods Acting as BOS
In Section 4.1 we talked about general contribution patterns and saw visible clusters corresponding to the input tokens’ part of speech. Now, let us go deeper and look in detail at one of the clusters. We choose the outlier punctuation cluster shown in yellow in Figure 8 (to the right) – this cluster corresponds to the first period in a text.777823 out of 826 datapoints.
Figure 9b shows the average importance of model components for examples in this cluster. We see that for these examples, the residual stream ignores all attention and FFN blocks in all the layers except for the first and last few: for most of the layers, contributions of all model components are near zero. When we look at the information flow graphs for these examples, we see that, even for a rather small threshold $\tau=0.01$, bottom-to-top processing happens largely via “untouched” residual connection (Figure 10). In Appendix C we show that up to the last three layers, this residual stream takes the role of the BOS token and future tokens treat this residual stream as such. Concurrent work (Cancedda, 2024) suggests that early FFN’s update into a particular subspace is responsible for this behavior. While for some examples periods acting as BOS might be reasonable, this also happens in cases where the period does not have an end-of-sentence role. For example, Figure 10 (right): while in a sentence For the second game in a row, St. Thomas ... the first period does not have the end-of-sentence meaning, the residual stream is still acting in the same manner. In future work, it might be valuable to explore whether this behavior might cause incorrect generation behavior.
@@ 5 Domain-Specific Model Components
We just saw that some of attention head functions are universally important across all predictions. Now, let us see whether some model components are not important universally but are specialized for specific types of data.
Setting.
We run Llama 2-7B on 1000 sentences from different domains and evaluate the average importance of model components within each domain. Specifically, we consider (i) C4 (Raffel et al., 2020), (ii) different languages: English, Russian, Italian, Spanish FloRes-200 devtest sets (NLLB et al., 2022; Goyal et al., 2022), (iii) code data from CodeParrot [^8] https://huggingface.co/datasets/codeparrot/github-code, (iv) addition/subtraction data (Stolfo et al., 2023).
@@ 5.1 Components are Specialized
Figure 11 shows the results: generic importance of attention heads and FFN blocks as well as their importance for specific datasets.
@@
From Figure 11 (left) we see that generally unimportant attention heads become highly relevant for specific tasks. For example, the most important heads for addition are among the lowest scoring heads when looking at C4 (Figure 11 (left), large yellow blobs to the left). Another interesting observation is that domain-specific heads are different across domains: important heads for addition, code, and non-English are not the same heads. Overall, we can conclude that model components are largely specialized.
@@
Figure 11 (right) shows more fine-grained results where instead of different domains, we look at tasks within the same narrow domain: addition vs subtraction. While the important heads for addition and subtraction largely intersect, we see several heads that are active only for one task and not the other (bright blue and yellow blobs). This means that this fine-grained specialization might be responsible for “reasoning” inside the model and not just domain-specific processing; future work may validate this further. Interestingly, we did not find similar specialization between languages and only find non-English heads – probably, portions of training data in these languages were not large enough to have dedicated language-specific heads.
Finally, we can also see the difference in the importance of the feed-forward blocks with respect to each domain. For example, the last FFN layer is much less relevant for non-English than other domains (Figure 11, left), the importance of FFN blocks for addition and subtraction falls to zero in some layers (Figure 11, right). In future work, one might conduct a more fine-grained analysis of the importance of feed-forward layers by looking at the relevance and functions of individual neurons (along the lines of Geva et al. (2021); Dai et al. (2022); Voita et al. (2023) but taking into account neuron importance for specific prediction).
5.2 Specialized Heads Output
Topic-Related Concepts
@@ 5.2 Specialized Heads Output
As we illustrated in Figure 4, $W^{h}_{OV}$ transforms representations from each of the residual streams [^9] After applying layer normalization first. into vectors that are added to the current residual stream. To understand what kind of information is embedded in this transformation, we use Singular Value Decomposition (SVD).
@@
For some of the heads specific to code and non-English inputs we saw in Figure 11, Figure 12 shows the top 10 tokens that come from the described above projection. We see that code-specific heads promote tokens related to coding and technology in a more general sense: e.g., tokens related to Apple, like iOS, Xcode, iPhone, etc. Heads most active for non-English promote tokens in multiple languages, avoiding English ones. Also, we see tokens related to cities (Sydney), counties (Malay), and currencies (£). Overall, we looked at the functions of attention heads from two perspectives: (i) when a head is active, and (ii) how it updates the residual stream. While a similar kind of analysis was done before for neurons (Voita et al., 2023), our method made it possible to talk about entire model components being active for a prediction or not.
@@ 6 Related Work
Earlier works evaluating the importance of model components include Bau et al. (2019), who identified important neurons in NMT models shared across models, and Voita et al. (2019) who looked at the relevance of entire attention heads in the Transformer from two perspectives: attribution and pruning. More recent research focuses more on task-specific rather than overall importance. They aim to find subparts of LMs responsible for these tasks and rely on activation patching methodology, firstly introduced to analyze biases in LMs by Vig et al. (2020), and used across several others (Heimersheim & Janiak, 2023). Initial works discovered circuits for different tasks of GPT2, like IOI (Wang et al., 2023) and greater-than (Hanna et al., 2023). Patching has been also used to locate factual knowledge in LMs (Meng et al., 2022), and to discover task vectors behind in-context learning capabilities (Hendel et al., 2023; Todd et al., 2023). Our work can be related to the line of research developing methods to automatize the discovery of circuits implemented by LMs (Syed et al., 2023; Conmy et al., 2023).
@@ 8 Acknowledgments
Appendix A Linearizing the Layer Normalization
Given an input representation ${\bm{x}}_{j}$, the layernorm computes
$\text{LN}({\bm{x}}_{j})=\frac{{\bm{x}}_{j}-\mu({\bm{x}}_{j})}{\sigma({\bm{x}}_{j})}\odot\mathbf{\gamma}+\mathbf{\beta}$ (5)
with $\mu$ and $\sigma$ obtaining the mean and standard deviation, and $\gamma\in\mathbb{R}^{d}$ and $\beta\in\mathbb{R}^{d}$ refer to learned element-wise transformation and bias respectively. Considering $\sigma({\bm{x}}_{j})$ as a constant, LN can be treated as a constant affine transformation:
$\text{LN}({\bm{x}}_{j})={\bm{x}}_{j}L+\mathbf{\beta}$ (6)
@@ Appendix A Linearizing the Layer Normalization
${\mathbf{L}:=\frac{1}{\sigma({\bm{x}}_{j})}\left[\begin{array}[]{cccc}\gamma_{1}&0&\cdots&0\\ 0&\gamma_{2}&\cdots&0\\ \cdots&\cdots&\cdots&\cdots\\ 0&0&\cdots&\gamma_{n}\end{array}\right]\left[\begin{array}[]{cccc}\frac{n-1}{n}&-\frac{1}{n}&\cdots&-\frac{1}{n}\\ -\frac{1}{n}&\frac{n-1}{n}&\cdots&-\frac{1}{n}\\ \cdots&\cdots&\cdots&\cdots\\ -\frac{1}{n}&-\frac{1}{n}&\cdots&\frac{n-1}{n}\end{array}\right]}$
The linear map on the right subtracts the mean to the input vector, ${\bm{x}}_{j}^{\prime}={\bm{x}}_{j}-\mu({\bm{x}}_{j})$. The left matrix performs the hadamard product with the layer normalization weights (${\bm{x}}_{j}^{\prime}\odot\gamma$).
Appendix B Examples of Routes
@@
Appendix C Period Acting as BOS

### v2

@@ 3.2 Greater-than
For the overall contributions in the greater-than task (Figure 6, left), we also see that (i) task-specific heads discovered via patching are among the most important heads in the information flow routes, and (ii) many other heads are important. Let us now look at the differences with the contrastive templates.
@@ 3.3 OPT-125m vs GPT2-Small
Additionally, we experiment with OPT-125m which has the same number of layers and heads than GPT2-small. We find that the information flow routes for IOI and greater-than tasks are similar for OPT-125M and GPT2-small (Appendix C).
@@ 3.4 Hundred Times Faster than Patching
Obtaining information flow routes involves a two-step process. First, we run a forward pass and cache internal activations. Then, we obtain edge importances to build the subgraph as depicted in Algorithm 3. For comparison purposes, we contrast our approach with the ACDC algorithm (Conmy et al., 2023), an automated circuit discovery algorithm based on patching. According to their findings, on the IOI task with a batch of 50 examples, it requires 8 minutes on a single GPU to discover the circuit. In contrast, our method accomplishes this task in 5 seconds, around 100x in time reduction.
@@ 4 General Experiments
Information flows inside the network via mechanisms implemented in the model. In this section, we look at general patterns in component importances, and try to understand if important mechanisms depend on the type of tokens processed. Using the full information flow graph (Figure 2) of Llama 2-7B, we look at all the immediate connections with each residual stream. We then record the importance values of all the sub-edges corresponding to individual attention heads (i.e., $e_{pos,j}^{h}$), as well as FFN blocks. We use a subset of 1000 sentences from the C4 dataset (Raffel et al., 2020).
@@ 4.1 Component Importance for POS
First, let us see how component importance depends on parts of speech (POS). For this, we pack per-prediction component importances into vectors [^5] Each vector corresponding to the $pos$-th position is defined as $\left(\sum_{j}e^{1,1}_{pos,j},\sum_{j}e^{1,2}_{pos,j},\ldots,\sum_{j}e^{L,H}_{pos,j},e^{\mbox{ffn}_{1}}_{pos},\ldots,e^{\mbox{ffn}_{L}}_{pos}\right)$. and apply t-SNE (van der Maaten & Hinton, 2008). Figure 8 shows the resulting projection with datapoints colored according to either input or next token [^6] Here, we take the next token from the dataset and not the one generated by the model. While this adds some noise to the results, we expect at least parts of speech of the reference and the predicted tokens to be similar. part of speech. [^7] We obtain POS tags with NLTK (universal tagset, Petrov et al. (2012)) and assign a tag to all subwords of each word.
@@
Diving deeper, Figure 8c shows that contribution patterns strongly depend on whether the current token is the first or a later subword of some word. Clearly, there are some model components specialized for first or later subwords – we confirm this later in Section 4.3. Note also that Figure 8c explains the two distinct clusters for numbers we saw in Figure 8a (purple): they separate number tokens into first and later digits.
@@
Finally, Figure 8b shows the same datapoint colored by the part of speech of the next token. Comparing this to Figure 8a, we see that contribution patterns depend more on input tokens rather than output tokens. This might be because in the lower part of the network, a model processes inputs in a generic manner, while the higher network part (that prepares to generate output) is more fine-grained and depends on more attributes than part of speech.
@@ 4.3 Positional and Subword Merging Heads
Some of our observations above hint at certain attention head functions that might be overall important. First, in Section 3.1 we showed that previous tokens heads are generally important in the IOI task, both for target examples and the contrastive baseline – this attention head function might be important in general. In Section 4.1 we noticed a clear difference between component contributions for first and later subwords – this suggests that some model components might be responsible for that.
In what follows, we look at two attention head functions found earlier for machine translation, positional and subword merging heads, and see whether they are generally important. Differently from previous work, we define a head function based on token contributions within the head and not attention weights as done previously, since attention weights might not reflect influences properly (Bastings & Filippova, 2020; Kobayashi et al., 2020, among others).
@@
Putting together our first-vs-later subword observations in Section 4.1 and previous work, we might expect our model to have subword merging heads found for machine translation encoders (Correia et al., 2019). We refer to a subword merging head if later subwords take information from previous subwords of the same word but the head is not previous token head (see Section B.2 for a formal definition). We see subword merging heads in the bottom part of the network (Figure 9a, red), and are among the most important heads in the corresponding layers. Notably, previous work did not study their general relevance (Correia et al., 2019). Note that these subword merging heads are important for later subwords and not important otherwise – this explains the clusters we saw in Figure 8c.
We would like to highlight that for LMs, we are the first to talk about general importance of attention heads and to find that some of the most important heads are previous token and subword merging heads. Future work might explore the functions of other important heads in the model.
@@ 4.4 Peculiar Information Flow Patterns, or Periods Acting as BOS
In Section 4.1 we talked about general contribution patterns and saw visible clusters corresponding to the input tokens’ part of speech. Now, let us go deeper and look in detail at one of the clusters. We choose the outlier punctuation cluster shown in yellow in Figure 8 (to the right) – this cluster corresponds to the first period in a text.888823 out of 826 datapoints.
Figure 9b shows the average importance of model components for examples in this cluster. We see that for these examples, the residual stream ignores all attention and FFN blocks in all the layers except for the first and last few: for most of the layers, contributions of all model components are near zero. When we look at the information flow graphs for these examples, we see that, even for a rather small threshold $\tau=0.01$, bottom-to-top processing happens largely via “untouched” residual connection (Figure 10). In Appendix D we show that up to the last three layers, this residual stream takes the role of the BOS token and future tokens treat this residual stream as such. Concurrent work (Cancedda, 2024) suggests that early FFN’s update into a particular subspace is responsible for this behavior. While for some examples periods acting as BOS might be reasonable, this also happens in cases where the period does not have an end-of-sentence role. For example, Figure 10 (right): while in a sentence For the second game in a row, St. Thomas ... the first period does not have the end-of-sentence meaning, the residual stream is still acting in the same manner. In future work, it might be valuable to explore whether this behavior might cause incorrect generation behavior.
@@ 5 Domain-Specific Model Components
We have seen that some attention heads’ roles are generally important across predictions. Now, we analyze if some model components are instead specialized in specific domains. We evaluate the average importance of Llama 2-7B’s components on 1000 sentences from different domains. Specifically, we consider (i) C4 (Raffel et al., 2020), (ii) English, Russian, Italian, Spanish FloRes-200 devtest sets (NLLB et al., 2022; Goyal et al., 2022), (iii) code data from CodeParrot [^9] https://huggingface.co/datasets/codeparrot/github-code, (iv) addition/subtraction data (Stolfo et al., 2023).
@@ 5.1 Components are Specialized
Figure 11 shows the importance of attention heads and FFN blocks in general and specific datasets.
@@
We see that generally unimportant attention heads become highly relevant for specific tasks. For example, the most important heads for addition are among the lowest scoring heads when looking at C4 (Figure 11 (left), large yellow blobs to the left). Another interesting observation is that domain-specific heads are different across domains: important heads for addition, code, and non-English are not the same heads. Overall, we can see that model components are largely specialized.
@@
Figure 11 (right) shows more fine-grained results where instead of different domains, we look at tasks within the same narrow domain: addition vs subtraction. While the important heads for addition and subtraction largely intersect, we see several heads that are active only for one task and not the other (bright blue and yellow blobs). This suggests that this fine-grained specialization might be responsible for “reasoning” inside the model and not just domain-specific processing; future work may validate this further. Interestingly, we did not find similar specialization between languages and only find non-English heads – probably, portions of training data in these languages were not large enough to have dedicated language-specific heads.
Finally, we can also see the difference in the importance of the feed-forward blocks with respect to each domain. For example, the last FFN layer is much less relevant for non-English than other domains (Figure 11, left), and the importance of FFN blocks for addition and subtraction falls to zero in some layers (Figure 11, right). In future work, one might conduct a more fine-grained analysis of the importance of feed-forward layers by looking at the relevance and functions of individual neurons (along the lines of (Geva et al., 2021; Dai et al., 2022; Voita et al., 2023) but taking into account neuron importance for specific predictions).
5.2 Specialized Heads Output Topic-Related Concepts
@@
For some of the heads specific to code and non-English inputs we saw in Figure 11, in Figure 12 we show the top 10 tokens that come from the described above projection. We see that code-specific heads promote tokens related to coding and technology in a more general sense: e.g., tokens related to Apple, like iOS, Xcode, iPhone, etc. Heads most active for non-English promote tokens in multiple languages, avoiding English ones. Also, we see tokens related to cities (Sydney), countries (Indones), and currencies (£). Overall, we looked at the functions of attention heads from two perspectives: (i) when a head is active, and (ii) how it updates the residual stream, and found they are consistent. While a similar kind of analysis was done before for neurons (Voita et al., 2023), our method made it possible to talk about entire model components being active/non-active for a prediction.
@@ 6 Related Work
Earlier works evaluating the importance of model components include Bau et al. (2019), who identified important neurons in NMT models shared across models, and Voita et al. (2019) who looked at the relevance of entire attention heads in the Transformer from two perspectives: attribution and pruning. More recent research focuses more on task-specific rather than overall importance. They aim to find subparts of LMs responsible for these tasks and rely on activation patching methodology, firstly introduced to analyze biases in LMs by Vig et al. (2020), and used across several others (Heimersheim & Janiak, 2023). Initial works discovered circuits for different tasks of GPT2, like IOI (Wang et al., 2023) and greater-than (Hanna et al., 2023). Patching has been also used to locate factual knowledge in LMs (Meng et al., 2022), and to discover task vectors behind in-context learning capabilities (Hendel et al., 2023; Todd et al., 2023). Our work can be related to the concurrent line of research developing methods to approximate activation patching (Syed et al., 2023; Kramár et al., 2024; Hanna et al., 2024).
@@ 8 Acknowledgments
Appendix A Background on Activation Patching
Activation patching (Vig et al., 2020; Meng et al., 2022; Geiger et al., 2020; Wang et al., 2023) refers to intervening some internal activation (intermediate representation) computed by a model component $c$ (attention head, feedforward network) in the forward pass ($f^{c}(x)$) with ‘base’ input $x$. The patched activation is taken from a forward pass $f^{c}(\tilde{x})$ on a ‘source’ input $\tilde{x}$. We can express this intervention using the do-operator (Pearl, 2009) as $f(x|\text{do}(f^{c}(x)=f^{c}(\tilde{x})))$. Upon intervention, the forward pass continues and the model output is compared with the prediction with the ‘base’ input, e.g. by measuring $f(x)-f(x|\text{do}(f^{c}(x)=f^{c}(\tilde{x})))$.
Identifying subnetworks (circuits) through activation patching has several shortcomings:
  • It requires large human efforts to create the input base templates ($x$) and contrastive source examples ($\tilde{x}$) for the specific task to study, thus results vary depending on the choice of the contrastive template. Additionally, analyses are constrained to those templates, preventing from studying models on more general types of predictions.
  • For each prediction, one needs to patch every edge (or node) in the computational graph, which becomes impractical when studying large language models.
  • It has been shown that downstream components can compensate for the ablation as a form of self-repair (McGrath et al., 2023; Rushing & Nanda, 2024), which interferes with the analysis.
In contrast, our method doesn’t require specific templates with contrastive examples which allows us to study specific tasks and more general behaviors, computes the information flow routes graph in a single forward pass, and it’s not affected by self-repair issues, since we make no interventions.
Appendix B Details about the Information Flow Routes
B.1 Linearizing the Layer Normalization
Given an input representation ${\bm{x}}$, the layernorm computes
$\text{LN}({\bm{x}})=\frac{{\bm{x}}-\mu({\bm{x}})}{\sigma({\bm{x}})}\odot\mathbf{\gamma}+\mathbf{\beta}$ (6)
with $\mu$ and $\sigma$ obtaining the mean and standard deviation, and $\gamma\in\mathbb{R}^{d}$ and $\beta\in\mathbb{R}^{d}$ refer to learned element-wise transformation and bias respectively. Considering $\sigma({\bm{x}})$ as a constant, LN can be treated as a constant affine transformation:
$\text{LN}({\bm{x}})={\bm{x}}L+\mathbf{\beta}$ (7)
@@ B.1 Linearizing the Layer Normalization
${\mathbf{L}:=\frac{1}{\sigma({\bm{x}})}\left[\begin{array}[]{cccc}\gamma_{1}&0&\cdots&0\\ 0&\gamma_{2}&\cdots&0\\ \cdots&\cdots&\cdots&\cdots\\ 0&0&\cdots&\gamma_{n}\end{array}\right]\left[\begin{array}[]{cccc}\frac{n-1}{n}&-\frac{1}{n}&\cdots&-\frac{1}{n}\\ -\frac{1}{n}&\frac{n-1}{n}&\cdots&-\frac{1}{n}\\ \cdots&\cdots&\cdots&\cdots\\ -\frac{1}{n}&-\frac{1}{n}&\cdots&\frac{n-1}{n}\end{array}\right]}$
The linear map on the right subtracts the mean to the input vector, ${\bm{x}}^{\prime}={\bm{x}}-\mu({\bm{x}})$. The left matrix performs the hadamard product with the layer normalization weights (${\bm{x}}^{\prime}\odot\gamma$).
B.2 Subword Merging Head definition
[…]
Appendix C Examples of Routes
@@
Appendix D Period Acting as BOS